moles of drink mix powder needed to make 3 solutions of 0.2 M,0.5 M and 0.8 M. Assume molar mass of powder is 180.156g/mol
1
Expert's answer
2019-01-08T03:05:06-0500

Answer

To make 0.2M solution one needs to take 0.2 mol of powder, put it into volumetric flask and add water until it reaches the line 1000 ml.

0.2 mole of powder has mass m= M*n= 180.156g/mol * 0.2 mol= 36.031 g.

To make 0.5M solution one needs to take 0.5 mol of powder, put it into volumetric flask and add water until it reaches the line 1000 ml.

0.5 mol of powder has mass m= M*n = 180.156 g/mol * 0.5 mol = 90.078 g

To make 0.8M solution one needs to take 0.8 mol of powder, put it into volumetric flask and add water until it reaches the line 1000 ml.

0.8 mol of powder has mass m=M*n = 180.156 g/mol * 0.8 mol = 144.125 g.